What quantity, x, of a 30% acid solution must be mixed with a 35% acid solution to produce 5000mL of a 31.5% solution?

Answer by lwsshak3(11628)

What quantity, x, of a 30% acid solution must be mixed with a 35% acid solution to produce 5000mL of a 31.5% solution?
***
x=amt of 30% solution to mix
5000-x=amt of 35% solution to mix
..
30%x+35%(5000-x)=31.5%*5000
.30x+.35(5000-x)=.315*5000
.30x+1750-.35x=1575
.05x=175
x=3500
5000-x=1500
amt of 30% solution to mix=3500 mL
amt of 35% solution to mix=1500 mL